OpenStudy (anonymous):

Find the distance between the points (4,-3) and (-6,2). HELP ME PLEASE

OpenStudy (ghazi):

\[\sqrt{(y2-y1)^2+(x2-x1)}\] use this

OpenStudy (anonymous):

Use distance formula: \[Distance = \sqrt{(x_2 - x_1)^2 + (y_2 -y_1)^2}\]

OpenStudy (ghazi):

\[\sqrt{(2+3)^2+(-10)^2}\]

OpenStudy (anonymous):

Sooo 25+ 100 then right?

OpenStudy (anonymous):

Go ahead..

OpenStudy (anonymous):

Sq Rt off 125 is idk haha

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

whats the answer i dont have a calculator handy

OpenStudy (ghazi):

\[\sqrt{125}= 5\sqrt{5}\]

OpenStudy (ghazi):

hope that helps you

OpenStudy (anonymous):

ok thanksss xD

OpenStudy (ghazi):

:)

OpenStudy (anonymous):

How do i find the mind point between (-1,4) and (-2,-9)?

OpenStudy (ghazi):

use mid point formula

OpenStudy (anonymous):

\[\sqrt{125} = \sqrt{\color{green}{\underline{5 \times 5}} \times5 } = 5 \sqrt{5}\]

OpenStudy (anonymous):

which is :)?

OpenStudy (anonymous):

Add the x coordinates and then divide them by 2..

OpenStudy (ghazi):

(-1+2)/2 and (4-9)/2

OpenStudy (anonymous):

Similarly do for y coordinates...

OpenStudy (anonymous):

so 0.5 and -2.5?

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

But that isnt an answer :(

OpenStudy (ghazi):

then it must be (1/2, -5/2)

OpenStudy (ghazi):

sorry x coordinate is -3/2

OpenStudy (anonymous):

the answers are (1.5,-5.5) (-2.5,2.5) (-1.5,-2.5) (1,6)

OpenStudy (anonymous):

well those are the options

OpenStudy (ghazi):

C

OpenStudy (anonymous):

M is the midpoint of segment AB. Find the coordinates of B given A(-5,1) and M(-2,4)

OpenStudy (anonymous):

Help help :)

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

add it and divide it by 2

OpenStudy (anonymous):

So just add -5+1? and the other one too ?

OpenStudy (ghazi):

(-7/2)(5/2)

OpenStudy (ghazi):

(-5-2)/2 and (4+1)/2

OpenStudy (anonymous):

oHh ohh ok so -6,4.5?

OpenStudy (ghazi):

no...(-3.5, 2.5)

OpenStudy (anonymous):

Oh sorry i did it wrong -__-

OpenStudy (anonymous):

In which region of angle CAB does point B lie?

OpenStudy (anonymous):